# Machine Learning Interatomic Potentials: library for efficient training, model development and simulation of molecular systems

### v1

@@ 4.1 Curated SPICE2 Dataset
We currently provide access to three pre-trained models, all trained on a curated second version of the SPICE2 dataset [56, 57]. The SPICE2 dataset was chosen for its diversity, both in chemical and conformational space, comprising approximately two million structures computed with DFT at the $\omega$B97M-D3(BJ)/def2-TZVPPD level of approximation. The dataset is labeled and subdivided into a collection of subsets; details of the training and validation sets per subset are provided in Table 1.
@@ 4.3 Model benchmarking
We evaluate the performance of the three pre-trained models, MACE-large, NequIP and VisNet, as well as the MACE-medium model (with hyperparameters aligned to MACE-OFF medium in [44]), on the SPICE2 validation set used during training. Each model was assessed with two standard error metrics: the root mean square error (RMSE) in predicted energies per atom (meV/atom) and the RMSE in atomic forces (meV/Å). Validation was conducted across seven subsets of SPICE2, including isolated and solvated small molecules, as described in Table 1.
Figure 2: Validation set root mean square errors (RMSE) for energy per atom (meV/atom) and atomic forces (meV/Å) across seven molecular subsets in the SPICE2 dataset. The three pretrained models— MACE-large, VisNet, and NequIP— as well as MACE-medium (following the MACE-OFF [44] hyperparameters) are evaluated. The subsets include: PubChem, DES370K, amino acid ligands, dipeptides, monomers, solvated PubChem, and solvated amino acids. RMSE values reflect the deviation from DFT reference calculations. The number of structures per subset is detailed in Table 1.
Figure 2 presents a comparative summary of model performance across all subsets. NequIP achieves the lowest energy RMSE for all subsets, while Visnet outperforms in force RMSE. Across most models, lower force errors are observed in the DES370K, dipeptides, monomers and solvated amino acids subsets than in the PubChem and solvated PubChem subsets. While MACE-medium achieves lower energy RMSE than MACE-large in every subset, MACE-large, on average, outperforms MACE-medium in force RMSE. Users should be warned that while validation errors are relevant metrics to measure training performance, they are not sufficient to attest to a model’s ability to simulate correct physics.
@@ 5 Roadmap for library development
We aim to release several updates to the current version (v0.1.0) in the coming months. These will include:

### v2

@@ 4.1 Curated SPICE2 Dataset
[…]
@@ 4.2 Model training methodology
Finally, we also trained a model with a modified MACE architecture optimized for inference speed, which we present in detail in Appendix C.
@@ 4.3 Model benchmarking
We evaluate the performance of the three pre-trained models, MACE-large, NequIP and VisNet, as well as the MACE-medium model (with hyperparameters aligned to MACE-OFF medium in [44]) and the modified MACE model (see Appendix C), on the SPICE2 validation set used during training. Each model was assessed with two standard error metrics: the mean absolute error (MAE) in predicted energies per atom (meV/atom) and the MAE in atomic forces (meV/Å). Validation was conducted across seven subsets of SPICE2, including isolated and solvated small molecules, as described in Table 1.
Figure 2: Validation set mean absolute errors (MAE) for energy per atom (meV/atom) and atomic forces (meV/Å) across seven molecular subsets in the SPICE2 dataset. The three pretrained models— MACE-large, VisNet, and NequIP— as well as MACE-medium (following the MACE-OFF [44] hyperparameters) and our modified MACE model (see Appendix C) are evaluated. The subsets include: PubChem, DES370K, amino acid ligands, dipeptides, monomers, solvated PubChem, and solvated amino acids. MAE values reflect the deviation from DFT reference calculations. The number of structures per subset is detailed in Table 1.
Figure 2 presents a comparative summary of model performance across all subsets. NequIP achieves the lowest energy MAE for all subsets, while Visnet outperforms in force MAE. Across most models, lower force errors are observed in the DES370K, dipeptides, monomers and solvated amino acids subsets than in the PubChem and solvated PubChem subsets. While MACE-medium achieves lower energy RMSE than MACE-large in every subset, MACE-large, on average, outperforms MACE-medium in force RMSE. Users should be warned that while validation errors are relevant metrics to measure training performance, they are not sufficient to attest to a model’s ability to simulate correct physics.
@@
Modified MACE 2,203,504 1UAO 3.0 ms/step 6.4 ms/step n.a.
1ABT 26.6 ms/step 48.7 ms/step n.a.
@@ 5 Roadmap for library development
We aim to release several updates to the current version (v0.1.3) in the coming months. These will include:
@@ Appendix B Appendix B - SPICE1 training of MACE medium
    Appendix C Appendix C - Modified MACE model
    In this Appendix, we introduce a modified version of MACE that exhibits similar performances to the original MACE model while significantly reducing the associated inference time. A key component of MACE is the so-called Symmetric Contraction (SC), a costly node-wise operation that computes the tensor product of node features with themselves $\nu$ times, where $\nu$ is referred to as the correlation order. The choice of $\nu$ depends on a speed-accuracy trade-off, with a correlation order $\nu=3$ leading to a slower but more accurate model than $\nu=2$. Here we propose two simple yet effective modifications of the original MACE model that overcome this tradeoff, allowing to reach an accuracy close the one of our vanilla MACE model with $\nu=3$ at an inference speed similar to the one obtained for $\nu=2$.
    In the following we describe our modifications by referring to the equations of the original MACE paper and adopt the same notations. Our first modification consists in applying a gating to the node features of node $i$ using the scalar produced during the SC. More precisely, we first define the following gating weights
    $\alpha^{(t)}_{Z_{i}kL,\eta_{\nu}}=\sum_{\tilde{\eta}_{\nu}}W^{\tilde{\eta}_{\nu}}_{Z_{i}k\tilde{k},\eta_{\nu}}\mathbf{B}^{(t)}_{i,\tilde{\eta}_{\nu}\tilde{k}00}+b_{Z_{i}k,\eta_{\nu}}$ (1)
    where $\mathbf{B}^{(t)}_{i,\tilde{\eta}_{\nu}\tilde{k}00}$ are the scalar features at the output of the SC and $W^{\tilde{\eta}_{\nu}}_{Z_{i}k\tilde{k},\eta_{\nu}},b_{Z_{i}k,\eta_{\nu}}$ are specie-wise learnable mixing weights and biases that depends on node $i$ through its atomic specie $Z_{i}$. Then, we gate the node features with the weights $\alpha^{(t)}_{Z_{i}kL,\eta_{\nu}}$, which amount to replace Eq. (11) defining the node messages with the following one:
    $m^{(t)}_{i,kLM}=\sum_{\nu}\sum_{\eta_{\nu}}\alpha^{(t)}_{Z_{i}kL,\eta_{\nu}}W^{(t)}_{Z_{i}kL,\eta_{\nu}}\mathbf{B}^{(t)}_{i,\eta_{\nu}kLM}\,$. (2)
    We found that this modification accounted for most of the improvement in our model. Importantly, this modification increases the body-order of the node-features while avoiding an increase in the correlation order, at a computational cost negligible compared to that of SC.
    The second modification relates to the interaction of a node with its neighbors as encapsulated in Eq. (8) of the original MACE paper. We generalize this interaction term as to let it depend explicitly on the species of both the sender and receiver nodes. To do so, we first embed the specie of each node $i$ into a vector of learnable scalar features
    $s_{i}:=(a^{1}_{Z_{i}},\dots,a^{d}_{Z_{i}})^{T}\in\mathbb{R}^{d}\,$, (3)
    with $d=8$. Then, we build scalar feature vectors on each edge $(i,j)$ as
[…]
    where $||$ and $\circ$ respectively denote concatenation and Hadamard product. The features $f_{ij}$ are then processed through a MLP to produce weights
    $\beta^{(t)}_{ij,kl_{1}l_{2}l_{3}}:=\mathrm{MLP}(f_{ij})_{kl_{1}l_{2}l_{3}}\,.$ (5)
    At last, Eq. (8) of the original MACE paper is to be replaced by the following equation:
    $A^{(t)}_{i,kl_{3}m_{3}}=\sum_{l_{1}m_{1},l_{2}m_{2}}C^{l_{3}m_{3}}_{l_{1}m_{1},l_{2}m_{2}}\sum_{j\in\mathcal{N}(i)}\beta^{(t)}_{ij,kl_{1}l_{2}l_{3}}R^{(t)}_{kl_{1}l_{2}l_{3}}(r_{ji})Y^{m_{1}}_{l_{1}}(\mathbf{\hat{r}}_{ji})\sum_{\tilde{k}}W^{(t)}_{k\tilde{k}l_{2}}h^{(t)}_{j,\tilde{k}l_{2}m_{2}}\,$. (6)